# Hedging Hydrogen: Planning and Contracting Under Uncertainty for a Green Hydrogen Producer

### v2

@@ Abstract
Green hydrogen production by water electrolysis using renewable electricity is considered essential for decarbonisation of certain sectors of the global economy, however development of the industry is lagging behind expectations due to the perceived financial risk for individual projects. This risk stems from a number of uncertainties, including future hydrogen demand, variable renewable energy sources, and volatile energy market prices. The interaction of these uncertainties is complex, yet the analysis of hydrogen projects is often carried out using simplified modelling. In this study, we define a set of planning methods (planning policies) in order to compare the effectiveness of different modelling approaches. We propose a 2-stage market-focused stochastic program to represent a hydrogen producer supplying an industrial customer through a hydrogen offtake contract (a Hydrogen Purchase Agreement, or HPA). The model can be used to obtain equipment sizing decisions, as well as energy hedging decisions using Power Purchase Agreements (PPA’s) and power futures. We find that for some HPA contract types, failure to use stochastic modelling can result in planning decisions that will perform 30% worse during scenario stress-testing for the same project. This could lead to some projects being incorrectly discarded. The results also show the importance of negotiating demand-side uncertainty in HPA contract negotiations.
@@ 1.1 Context
As part of the energy transition and the fight against climate change, expanding the use of green hydrogen produced from water and renewable energy through electrolysis is considered to be a promising method of reducing emissions in many hard-to-decarbonise sectors of the economy. Aside from decarbonising current use cases of hydrogen such as fertiliser production and oil refining, green hydrogen may be used in the future to replace fossil fuels in logistics, in the creation of sustainable shipping and aviation fuels, as well as a chemical agent or feedstock in certain industrial processes [1].
Development of the low-carbon hydrogen economy [1] requires investment in both supply and demand. Production necessitates investment in expensive electrolysers, and use of hydrogen on the demand side often requires re-tooling or new equipment. There is also considerable uncertainty in production costs for green hydrogen, as they are closely related to electricity prices which are highly volatile and sensitive to geopolitical shocks.
Attracting the necessary financing to cover these high capital costs can be difficult given the wariness of lenders with respect to the aforementioned risks [2], as well as the high cost of capital and slow implementation of regulation [3]. As a result, green hydrogen project investment is lagging behind the pace and volume required to meet global climate ambitions [3, 4].
Green hydrogen producers require long-term confidence in demand for their product to make their projects bankable and attract the necessary financing to cover significant capital costs [5]. Especially given that markets for green hydrogen do not currently exist, the natural solution for this uncertainty is to find industrial customers willing to engage in bilateral offtake contracts, or Hydrogen Purchase Agreements (HPA’s) [6], wherein a future producer and a future consumer of hydrogen commit to exchanges of hydrogen over an extended period.
In order to sign up to a HPA, green hydrogen consumers require long-term confidence in competitive prices for green hydrogen to avoid “the risk of becoming locked into an expensive and potentially scarce energy carrier” [4]. Potential consumers must be reassured that their hydrogen purchase price will stay competitive with alternatives over the life of the contract even in adverse energy market conditions.
@@
The aim of this paper is therefore to investigate the relative effectiveness of different types of quantitative planning methods (or Planning Policies using the terminology of [7]) for decision making in green hydrogen production in the context of a HPA contract.
With the aim of reducing financial risk, green hydrogen producers have several risk mitigation mechanisms at their disposal which can be separated into the two categories of asset sizing, or energy hedging. Asset sizing decisions may include oversizing of certain flexible assets (such as hydrogen storage) to allow increased utilisation of renewable sources or to ride through market price peaks. Energy hedging decisions relate to the procurement of electricity in advance outside of the spot markets. The most commonly used hedging instruments to date being power futures (where electricity prices are locked in for block volumes with delivery at a future date) [8], and renewable Power Purchase Agreements (PPA’s) — non-standard contracts for which the most typical form (pay-as-produced) locks in a set price at which a set proportion of production of a given renewable energy park is bought by the consumer as it is generated [9].
Additional help may be available in the form of incentives for green hydrogen producers [10], such as subsidy auctions in Europe [11], and a production tax credit in the US [12]. Eligibility for these subsidy schemes is to require hourly time-correlation of hydrogen production with renewable sources [13]. As such, the benefit of achieving eligibility to receive these subsidies will influence electricity procurement decisions for individual projects.
The interplay between these sources of uncertainty and potential instruments for financial risk mitigation is far from simple, and yet in both industry and academia approximations are commonly used in order to simplify modelling and decision-making (which are discussed in the following section). This study draws from existing literature in three thematic areas: (1) techno-economic analysis of hydrogen electrolysis plants, (2) capacity expansion planning under uncertainty, and (3) electricity futures trading under uncertainty.
@@ 1.2 Literature Gaps and Paper Objectives
1.2.1 Technical Economic Analysis of Hydrogen Projects
Simulation-based feasibility studies for the production of green hydrogen via pay-as-produced PPA’s can be found in [14, 15]. Both studies are however performed in a deterministic setting and without optimising capacities. The feasibility study [16] does include capacity expansion and compares several energy procurement configurations, although also in a deterministic setting. In [17] a capacity expansion model is proposed for a hydrogen refuelling station, again without the presence of uncertainty. In [18] a capacity expansion model is used to perform a sensitivity analysis of hydrogen cost and emissions intensity with respect to average electricity costs and under various regulatory contexts in a deterministic setting. A review of system level hydrogen capacity expansion models is given in [19], while [20] and [21] examine the system level impacts of different hydrogen regulatory rules in the US and Europe. Anecdotal evidence suggests deterministic modelling is also widespread in industry for project planning and valuation.
@@ 1.2.2 Hydrogen Capacity Expansion Planning Under Uncertainty
Making capacity expansion decisions with the presence of uncertainty is a problem commonly addressed by stochastic optimisation [22]. In [23] a stochastic capacity expansion model is proposed for sizing of an electrolytic hydrogen system with uncertainty in both solar production and hydrogen equipment CAPEX, although neither demand uncertainty nor grid connection are included. In [24], electrolysis plant and HVDC link capacities are optimised at a system level with both spot price and renewable production uncertainty, however neither demand uncertainty nor energy hedging are included. Outside the hydrogen domain, [25] proposes a capacity expansion model in the context of an industrial customer with uncertain demand and energy prices, however renewable production and energy hedging are not included. In [26] a stochastic capacity expansion model with risk aversion is used to optimise the market performance of a portfolio of PPA’s, although without demand uncertainty, futures hedging, or storage.
@@ 1.2.3 Electricity Hedging Under Uncertainty
To the authors’ knowledge, there are no studies specifically addressing hedging in the context of hydrogen projects. In the energy industry more broadly, hedging decisions using heuristics such as predefined hedging ratios are common [8]. Some studies address futures hedging for electro-intensive industrial consumers, however energy hedging is often studied in an operational manner, and equipment capacity expansion is not included. Stochastic programming approaches for futures hedging for an energy intensive consumer are given in [27] and [28]. A more recent study applies a similar approach to the case of energy futures hedging for an air-separation unit [29]. A rare example where both equipment capacity investment and futures trading are combined is given in [30], however its agent-based modelling requires assumptions about maximum capacities available to different market players, which are not appropriate for a single price-taker interested in hedging their risk. One thesis that shall be considered in this study is that capacity sizing decisions affect optimal procurement decisions, and vice-versa, and so performing a comprehensive optimisation of both capacity planning and procurement strategy simultaneously is preferable.
Risk-Neutral Pricing and Speculation
@@
In these previous studies, an important modelling decision is how futures should be priced with respect to the spot market. The use of arbitrary prices (such as in [28]) may inadvertently make the purchase of futures attractive for their resale value alone — not for alleviating risk. This can lead to excessive speculation within the model. The term speculation is used in this article to refer to the procurement of energy with the express purpose of reselling that energy on the spot market for profit, rather than using it to fulfil the underlying demand process 111This action is often termed arbitrage in energy modelling literature [30], however this differs from strict definitions of arbitrage in finance theory [31]..
A commonly used approach is to adjust either the scenarios, or the hedging instrument prices, so that speculation has zero profitability on average, a state that is often referred to as risk-neutral pricing [32]. In [33] a model for dynamic futures hedging for a hydropower producer uses risk-neutral day-ahead scenarios produced from known futures prices. In [27, 34], the reverse is done: risk-neutral futures prices are calculated based on the expectation of day-ahead market simulations from spot price models.
1.2.4 No-Resale Formulations in Hedging Problems
If risk-neutral pricing is not used for a given hedging option, the decisions of a risk-neutral optimisation will almost certainly include an element of speculation. Excessive speculation may result in decisions such as the purchasing of energy futures or PPA’s up to the maximum allowable bounds when the expected value of the revenue from their resale is greater than the cost of procuring them.
@@ 1.2.4 No-Resale Formulations in Hedging Problems
In order to avoid this problem without using risk-neutral pricing, a common feature of these aforementioned energy hedging studies [27, 29, 30] is the limitation (or more often, complete suppression) of hedging energy resale on the spot market. Modelling formulations that limit hedging energy re-sale shall be referred to as no-resale formulations. These formulations represent a departure from reality, as excess energy can usually be sold on the spot market. Indeed in practice it must, as hedging instrument delivery profiles will not cover exactly the profile of the demand to be hedged, and so systematic buying the shortfall and selling the surplus on the spot market will occur even if 100% hedged on yearly volume. A thesis proposed in this paper is that no-resale formulations should not be used when deciding between different hedging options.

### v3

@@ Abstract
Green hydrogen production by water electrolysis using renewable electricity is considered essential for decarbonisation of certain sectors of the global economy, however development of the industry is lagging behind expectations due to the perceived financial risk for individual projects. This risk stems from a number of uncertainties, including future hydrogen demand, variable renewable energy sources, and volatile energy market prices. The interaction of these uncertainties is complex, yet the analysis of hydrogen projects is often carried out using simplified modelling that often omits uncertainty and/or energy hedging practices which are typical for intensive power consumers. In this study, we define a set of planning methods (planning policies) in order to compare the effectiveness of different modelling approaches. We propose a 2-stage market-focused stochastic program to represent a hydrogen producer supplying an industrial customer through a hydrogen offtake contract (a Hydrogen Purchase Agreement, or HPA). The model can be used to obtain equipment sizing decisions, as well as energy hedging decisions using Power Purchase Agreements (PPA’s) and power futures. We find that for some HPA contract types, failure to use stochastic modelling can lead to planning decisions that result in 30% higher production costs during scenario stress-testing for the same project. This could lead to some projects being discarded by developers, incorrectly deemed to be unviable due to cost projections being too high. The results also show the importance of HPA contract volumetric obligations in limiting demand uncertainty.
@@
An energy market-focused 2-stage stochastic model of an electrolytic hydrogen plant.
Co-optimisation of sizing and energy hedging for Hydrogen Purchase Agreements (HPA).
Deterministic planning methods may cause viable projects to be overlooked.
Fixed-demand HPAs lower CAPEX, while uncertain demand raises hydrogen sale prices.
@@ 1.1 Context
As part of the energy transition and the fight against climate change, expanding the use of green hydrogen produced from water and renewable energy through electrolysis is considered to be a promising method of reducing emissions in many hard-to-decarbonise sectors of the economy. Aside from decarbonising current use cases of hydrogen such as fertiliser production and oil refining, green hydrogen may be used in the future to replace fossil fuels in logistics, in the creation of sustainable shipping and aviation fuels, as well as a chemical agent or feedstock in certain industrial processes (InternationalEnergyAgency2019TheOpportunities).
Development of the low-carbon hydrogen economy (InternationalEnergyAgency2019TheOpportunities) requires investment in both supply and demand. Production necessitates investment in expensive electrolysers, and use of hydrogen on the demand side often requires re-tooling or new equipment. There is also considerable uncertainty in production costs for green hydrogen, as they are closely related to electricity prices which are highly volatile and sensitive to geopolitical shocks.
Attracting the necessary financing to cover these high capital costs can be difficult given the risk-aversion of lenders with respect to the aforementioned risks (GreenHydrogenOrganisation2022GreenProjects), as well as the high cost of capital and slow implementation of regulation (InternationalEnergyAgency2023Global2023). As a result, green hydrogen project investment is lagging behind the pace and volume required to meet global climate ambitions (InternationalEnergyAgency2023Global2023; Odenweller2024TheGap).
Green hydrogen producers require long-term confidence in demand for their product to make their projects bankable and attract the necessary financing to cover significant capital costs (Craen2023FinancingProject; Scholvin2025De-riskingAfrica). Especially given that markets for green hydrogen do not currently exist, the natural solution for this uncertainty is to find industrial customers willing to engage in bilateral offtake contracts, or Hydrogen Purchase Agreements (HPA’s), wherein a future producer and a future consumer of hydrogen commit to exchanges of hydrogen over an extended period (Craen2023FinancingProject; Scholvin2025De-riskingAfrica).
In order to sign up to an HPA, green hydrogen consumers require long-term confidence in competitive prices for green hydrogen to avoid “the risk of becoming locked into an expensive and potentially scarce energy carrier” (Odenweller2024TheGap). Potential consumers must be reassured that their hydrogen purchase price will stay competitive with alternatives over the life of the contract even in adverse energy market conditions.
@@
The aim of this paper is therefore to investigate the relative effectiveness of different types of quantitative planning methods (or Planning Policies using the terminology of Powell2022SequentialII) for decision making in green hydrogen production in the context of an HPA contract.
With the aim of reducing financial risk, green hydrogen producers have several risk mitigation mechanisms at their disposal which can be separated into the two categories of asset sizing, or energy hedging. Asset sizing decisions may include oversizing of certain flexible assets (such as hydrogen storage) to allow increased utilisation of renewable sources or to ride through market price peaks. Energy hedging decisions relate to the procurement of electricity in advance outside of the spot markets. The most commonly used hedging instruments to date being power futures (where electricity prices are locked in for block volumes with delivery at a future date) (Sanda2013SelectiveCompanies), although renewable Power Purchase Agreements (PPA’s) are becoming more and more common (Flottmann2024DerivativesMarket). PPA’s are non-standard contracts for which the most typical form, pay-as-produced, locks in a set price at which a set proportion of production of a given renewable energy park is bought by the consumer as it is generated (Mittler2023ReviewAnalysis).
Additional help may be available in the form of incentives for green hydrogen producers (Hoogsteyn2025InteractionsHydrogen), such as subsidy auctions in Europe (EuropeanCommission2023QuestionsHydrogen), and a production tax credit in the US (USDepartmentofEnergy2023AssessingCredit). Eligibility for these subsidy schemes is to require hourly time-correlation of hydrogen production with renewable sources (Guillotin2025HydrogenComparison). As such, the benefit of achieving eligibility to receive these subsidies will influence electricity procurement decisions for individual projects.
The interplay between these sources of uncertainty and potential instruments for financial risk mitigation is far from simple, and yet in both industry and academia approximations are commonly used in order to simplify modelling and decision-making (which are discussed in the literature review to follow). This study builds on literature in three thematic areas: (1) techno-economic analysis of hydrogen electrolysis projects, (2) capacity expansion planning under uncertainty, and (3) electricity hedging under uncertainty.
@@ 1.2 Literature Gaps and Paper Objectives
1.2.1 Techno-Economic Analysis of Hydrogen Projects
Many techno-economic feasibility and capacity expansion studies in the hydrogen domain can be found, however the majority are performed using deterministic modelling which inherently ignores uncertainty, and overlooks the risk-averse preferences of investors and project developers. The impact of these simplifications in the context of hydrogen modelling is yet to be quantified in a scientific study. In reviews of modelling tools available for energy system modelling (Haugen2023RepresentationReview), and specifically in the hydrogen domain (Reulein2025TheModels), it is noted that “most applied power system tools are deterministic” (Haugen2023RepresentationReview). Anecdotal evidence suggests that deterministic modelling is indeed common in industry for project planning and valuation.
Some examples of this in scientific literature include (AliKhan2021DesigningAustralia; Matute2023Techno-economicPPAs), in which simulation-based feasibility studies for green hydrogen production via pay-as-produced PPA’s is presented, although in a deterministic setting and without optimising capacities. A similarly deterministic model without capacity expansion is proposed in Koleva2021OptimalMarkets, in which the authors use iterative solutions of the model to improve electrolyser capacity solutions for different energy procurement configurations. Brandt2024CostFramework use a capacity expansion model to study hydrogen cost and emissions sensitivity to electricity prices, however the model used is again deterministic.
Several studies examine system-level impacts of hydrogen policies and incentives, although also in a deterministic setting (Ricks2023MinimizingStates; Shirizadeh2023Long-termHydrogen; Zeyen2024TemporalHydrogen; Vargas-Ferrer2025ComplyingChains; Hoogsteyn2025InteractionsHydrogen). As noted by Hoogsteyn2025InteractionsHydrogen in the limitations of their study, the assumption of perfect foresight (which is implicit to deterministic modelling) prevents deterministic models from taking into account risk-mitigating behaviour of investors and project developers.
@@ 1.2.2 Hydrogen Capacity Expansion Planning Under Uncertainty
There are scientific studies of capacity expansion under uncertainty in the hydrogen domain and the energy domain more broadly (Roald2022PowerApplications), however no single study takes into account all the necessary sources of uncertainty, flexible assets, and especially hedging instruments, all of which are required for a practical solution of the problem addressed in this paper.
For example, the studies by DeWeerdt2023TheDecisions and Fabianek2024ACalifornia apply real options approaches to hydrogen generator retrofitting and to electrolyser installation respectively, however these studies examine the tipping points between different equipment investment decisions, and how these tipping points change over time — they do not consider hedging of a contracted demand.
Similar limitations can be seen in studies using stochastic optimisation approaches for hydrogen system planning. Coppitters2019Surrogate-assistedUncertainty propose a stochastic capacity expansion model for sizing of an electrolysis plant with uncertainty in both solar production and equipment CAPEX, although neither demand uncertainty nor grid connection is included. Li2022CoordinatedUtilization perform a stochastic optimisation of electrolyser and HVDC capacities under spot price and renewable production uncertainty, though demand uncertainty and hedging are not considered.
Stochastic capacity expansion studies outside the hydrogen domain also provide some but not all necessary components. Mitra2014OptimalModeling propose a capacity expansion model for an industrial consumer with uncertain demand and energy prices, however variable renewable production and energy hedging are not included. A price-taker stochastic model for risk-averse generation investment is given in (Pisciella2016AExpansion), although it does not include demand and renewable production uncertainty, storage, or hedging instruments. Gal2017FuelMarket study gas generator capacity decisions, comparing with and without rule-based options hedging. The authors note that increasing price volatility leads to increased equipment capacity investment, however their study does not include risk-aversion, hydrogen production, or storage, and it uses a price-maker equilibrium model that is not appropriate for a single market participant interested in hedging their risk.
@@ 1.2.3 Electricity Hedging Under Uncertainty
There is a rich variety of hedging studies in the energy domain, however examples within the sub-domain of hydrogen production are rare. In addition, energy hedging is typically studied in an operational manner (i.e., with fixed physical asset sizes), and co-optimisation of equipment capacity is usually not included. This ignores the impact of equipment sizing on energy procurement decisions for flexible power-intensive consumers like hydrogen producers.
Energy hedging strategies can be defined as static (unchanging as new information is revealed), and dynamic (changing over time to adapt to new information, Dimoski2023DynamicRisks). Hedging volumes are often expressed in terms of the hedging ratio — the proportion of the expected future volume which is hedged. In the energy industry, hedging decisions using heuristics such as predefined static hedging ratios are common (Sanda2013SelectiveCompanies), which typically prescribe the purchase of hedging instruments incrementally over time to achieve a planned hedging ratio for an expected future volume.
There are few studies of energy hedging in the context of hydrogen production. A rare example is Wu2022ElectricityMarkets, who propose an operational stochastic program including futures hedging for a hydrogen recharging station, although capacity expansion is not included.
The hydrogen production hedging problem is however analogous to other studies relating to hedging for electricity-intensive consumers, although once again capacity expansion is usually not included. This is true for several studies which apply stochastic optimisation in an operational manner in this context (Carrion2007AConsumers; Conejo2010DecisionMarkets; Zhang2018Long-TermUncertainty). In an approach similar to that taken in this article, Gabrielli2022MitigatingOptimization present a stochastic hedging model with risk aversion to optimise a portfolio of PPA’s for demand fulfilment, however this study does not include hydrogen production, demand uncertainty, futures hedging, or storage.
Studies considering the problem of hedging for an electricity retailer also have useful parallels for the hydrogen planning problem studied in this paper, though these studies usually do not incorporate all the necessary elements. For example, (Boroumand2015HedgingRetailers; Ernstsen2017HedgingCase; Tegner2017Risk-minimisationConsumption) use similar methodologies to compare static hedging policies using risk metrics like Value at Risk (VAR) and Conditional Value at Risk (CVAR) for electricity retailers with and without means of generation, however these studies do not include capacity expansion, and variable renewable energy is not included.
Other works analyse hedging from a power producer’s perspective (Pineda2013UsingProducers; Pircalabu2017JointApproach; Tranberg2020ManagingAgreements; Thakur2023PricingContracts). While these studies provide relevant techniques with respect to pricing and scenario construction, they are essentially the inverse problem of the consumer-side problem addressed in this paper.
Some studies examine only the impact of hedging on profit and loss (P&L) with respect to an open position on the spot market, however this does not take into account the impact of demand profiles on optimal hedging strategy. One example of this is Alexander2013TheSpread which compares the effectiveness of several hedging policies for crude oil products. The authors note that simple hedging strategies can be effective in certain circumstances. In the electricity domain, a similar study is carried out by Hanly2018TheMarkets. In addition to the omission of demand profile effects, these P&L studies also omit renewables, capacity expansion, and storage, and are therefore of limited use for the hydrogen planning problem.
Co-optimisation of capacity expansion planning and procurement strategy is scarce in scientific literature, and the rare examples performing this use system-level equilibrium models which are not suitable for a single actor interested in hedging their risk. One example is the study by Kazempour2012StrategicMarkets, in which generation capacity investment and futures trading are optimised. However its agent-based modelling requires assumptions about maximum capacities available to different market players, which is not realistic for a single price-taker with limited information. A different approach is taken by Shu2023BeyondObligations, who use iterative solutions of a stochastic equilibrium model to optimise generator investment and different hedging behaviour under different risk-aversion levels and policy obligations. Futures, options, and PPA’s are included, however hydrogen production and storage are not. The same model is adjusted by FarhadBillimoria2024HedgingMarkets to include the possibility of battery storage, and it is then used to study changes in hedging costs in a progressively renewable grid. However, this iterative method lacks guaranteed optimality and the equilibrium model remains unsuitable for single-actor hedging.
1.3 Speculation and No-Resale Formulations in Hedging Problems
Some studies in the electricity hedging domain can also be further characterised as using a no-resale formulation, the effects of which have not yet been studied in literature (to the authors’ knowledge). This modelling concept shall be introduced in the following sections.
1.3.1 Risk-Neutral Pricing
@@
An important modelling decision is how electricity futures should be priced with respect to the spot market. The use of arbitrary or unrelated prices (such as in Wu2022ElectricityMarkets and some studies of Conejo2010DecisionMarkets) may inadvertently make the purchase of futures attractive for their resale value alone — not for alleviating risk. This can lead to excessive speculation within the model. The term speculation is used in this article to refer to the procurement of energy with the express purpose of reselling that energy on the spot market for profit, rather than using it to fulfil the underlying demand process111This action is often termed arbitrage in energy modelling literature (Conejo2010DecisionMarkets; Kazempour2012StrategicMarkets; Pineda2013UsingProducers), however this differs from strict definitions of arbitrage in finance theory (Delbaen2008TheArbitrage).. A commonly used approach is to adjust either the scenarios, or the hedging instrument prices, so that speculation has zero profitability on average, a state that is often referred to as risk-neutral pricing (Deschatre2021AApplications). Fleten2003ConstructingMarkets and Dimoski2023DynamicRisks use risk-neutral day-ahead scenarios produced from known futures prices. In (Carrion2007AConsumers; Boroumand2015HedgingRetailers; Tranberg2020ManagingAgreements), the reverse is done: risk-neutral futures prices are calculated based on the expectation of day-ahead market simulations from spot price models.
1.3.2 No-Resale Formulations
If risk-neutral pricing is not used for a given hedging option, the decisions of a risk-neutral optimisation will almost certainly include an element of speculation. As noted by Boroumand2015HedgingRetailers, not using risk-neutral futures prices would mean that the futures contract valuation method will become the main driver for the results.
@@ 1.3.2 No-Resale Formulations
In order to avoid this problem without using risk-neutral pricing, a common feature of these aforementioned studies (Carrion2007AConsumers; Zhang2018Long-TermUncertainty; Kazempour2012StrategicMarkets; Wu2022ElectricityMarkets; Matute2023Techno-economicPPAs; Gabrielli2022MitigatingOptimization) is the limitation (or more often, complete suppression) of hedging energy resale on the spot market. Modelling formulations that limit hedging energy resale shall be referred to as no-resale formulations). These formulations represent a departure from reality, as excess energy can usually be sold on the spot market. Indeed in practice it must, as hedging instrument delivery profiles will not cover exactly the profile of the demand to be hedged, and so systematic buying the shortfall and selling the surplus on the spot market will occur even if 100% hedged on yearly volume. A thesis proposed in this paper is that no-resale formulations should not be used when deciding between different hedging options.
1.3.3 Summary and Literature Gaps
Despite the fact that hydrogen projects involve a complex mix of flexible assets, subsidy structures, numerous sources of uncertainty, and are typically developed by risk-averse agents, to the authors’ knowledge no existing study presents a complete method for jointly optimising equipment investment and electricity hedging for these projects. Furthermore, the impact of common modelling practices such as deterministic modelling, risk neutrality, and no-resale formulations is not known, despite the prevalence of these methods in academia and in industry.